OpenStudy (anonymous):

a)write the expression 32 ^-3/5 in radical notation. b)Evaluate the radical expression.

OpenStudy (tkhunny):

\('5' \implies \sqrt[5]{Something}\) \("-3" \implies \dfrac{1}{Something^{3}}\)

OpenStudy (anonymous):

I think its this: \[\frac{ 1 }{ \sqrt[5]{32^{3}} }\] Not entirely sure

OpenStudy (anonymous):

how do I evaluate the expression ?

OpenStudy (anonymous):

take the fifth root of 32 first the fifth root of 32 is 2, because \(2^5=32\)

OpenStudy (tkhunny):

It is easier this way: \(\dfrac{1}{\left(\sqrt[5]{32}\right)^{3}}\)

OpenStudy (anonymous):

then take 2 and cube it. \(2^3=2\times 2\times 2=8\)